OpenStudy (anonymous):

write the expression as a complex number in standard form. (10-4i)(3+4i)

OpenStudy (mathstudent55):

Use FOIL to multiply together like we did yesterday: (10 - 4i)(3 + 4i) = 10 * 3 + 10 * 4i - 4i * 3 - 4i * 4i Now do all 4 multiplications

OpenStudy (anonymous):

10*3=30 10*4i=40i 4i*3=12i 4i*4i=16i

OpenStudy (mathstudent55):

Almost. Look at the last one: 10*3=30 10*4i=40i 4i*3=12i 4i*4i=16i^2 = 16 * (-1) = -16

OpenStudy (anonymous):

so -16 how do we put it all together

OpenStudy (mathstudent55):

10 * 3 + 10 * 4i - 4i * 3 - 4i * 4i = 30 + 40i - 12i - (-16)

OpenStudy (mathstudent55):

10 * 3 + 10 * 4i - 4i * 3 - 4i * 4i = 30 + 40i - 12i - (-16) = 30 + 28i + 16 = 46 + 28i

OpenStudy (anonymous):

5/1+2i

OpenStudy (mathstudent55):

You mean this, right? 

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

You need to multiply the numerator and denominator by the complex conjugate of the denominator.

OpenStudy (mathstudent55):

Can you write the complex conjugate of the denominator?

OpenStudy (anonymous):

1-2i

OpenStudy (mathstudent55):

Right. Now we need to multiply.

OpenStudy (mathstudent55):